OpenStudy (anonymous):

a tennis player uses a strobe lamp to measure the speed of his serves. The lamp flashes every 1/100 of a second. A camera records the ball moving 31cm between flashes. what is the speed of the ball in m/s?

OpenStudy (anonymous):

v = delta x / delta t delta t = 1/100 s delta x = 31 cm m/s -> 100cm : 1 m .31/(1/100) = 31m/s I'm pretty sure. I took ap physics b last year, this is pretty basic kinematics.

OpenStudy (anonymous):

thanks :) this was a little advanced for a 7th grader like me :(

OpenStudy (anonymous):

no problem :) i miss 7th grade physics :p